# Zeyu Shen1    Basileal Imana1    Tong Wu1    Chong Xiang2    Prateek Mittal1    Aleksandra Korolova1

### v1

@@ D.1.2 Generality of Weighted Approaches
  • In keyword aggregation in [61], instead of simple counts, one could accumulate the sum of weights of documents supporting each keyword. The filtering threshold $\mu$ could then be applied to these weighted sums.
  • In decoding aggregation in [61], the averaging of next-token probability vectors could become a weighted average, using weights derived from the documents supporting each prediction $v_{j}$.
@@
D.2 Limitations and Future Work
@@
The efficacy of our MIS-based approach is intrinsically linked to the NLI model’s ability to accurately detect contradiction. Although Theorem 3.1 accounts for imperfect NLI, the practical impact of more severe NLI inaccuracies, or NLI models that are themselves targeted by sophisticated adversarial examples, deserves more study.
Computational Cost.
@@
Although exact MIS is practical for the typical number of retrieved documents (e.g. $k\leq 20$), and our weighted sample and aggregate framework extends scalability, the “isolated answering” step for contradiction graph construction (Section 3.1) does add nontrivial computational latency. While we have demonstrated that this overhead is manageable and have also provided practical speed-up tips, it is a factor to consider. Exploring more efficient methods for contradiction detection can be an interesting future direction.
Heuristic Choices of Parameters and Algorithmic Designs.
@@
Our proposed framework incorporates several design choices and parameter settings. For example, Algorithm 1 selects the MIS with the smallest lexicographic order. In our evaluations, we focused on specific configurations such as $m=2$, $T=20$, and using exponentially decaying weights with $\gamma=0.9$. While these configurations have demonstrated strong performance, and Appendix C.4 provides some analysis of how certain parameter choices affect performance, many other reasonable design choices remain interesting to explore. For instance, exploring the use of a maximum weighted independent set could offer a more direct integration of cardinal reliability scores into the MIS selection process itself. Another promising heuristic worth investigating involves applying Algorithm 1 recursively to filter each sampled document set $\mathcal{S}_{t}$ prior to generating intermediate answers in Algorithm 2 (Line 4), which might further bolster the reliability of the final aggregated response.
Reliance on LLM-as-a-Judge.
@@
Our empirical evaluations rely on GPT-4o as an LLM-judge for answer correctness and quality. While a common practice, LLM-based evaluation may have inherent biases and may not fully capture all nuances of human assessment.
Exploration of Diverse Adaptive Attack Strategies.
@@
Our current work evaluates robustness against several attack types, including corpus poisoning attack, prompt injection attack, and a specific adaptive attack scenario (as detailed in Appendix C.6). However, the landscape of adversarial tactics is continually evolving. To more comprehensively ascertain the resilience of ReliabilityRAG, future work should explore a wider array of sophisticated adaptive attacks. Adversaries with deeper knowledge of the defense mechanism might devise strategies not covered in our present evaluations. A thorough investigation of such advanced adaptive threats would further solidify the understanding of our method’s robustness boundaries and is a valuable direction for continued research.
Scope of Evaluation Benchmarks
@@
While our empirical evaluations utilize established datasets such as RQA, NQ, TQA, and the Biography generation dataset, which are common benchmarks in RAG research, it is important to acknowledge a potential limitation shared across much of the current literature. The characteristics and complexities of queries and documents encountered in these datasets may not fully encapsulate the diverse and dynamic nature of real-world web search combined with RAG systems. Consequently, while our results demonstrate significant robustness and utility, performance in live, large-scale commercial search + RAG environments might present additional, unforeseen challenges. This gap between academic benchmarks and real-world deployment scenarios is a broader issue faced by the research community.
@@
Our implementation uses an “I don’t know” response from an LLM to filter irrelevant documents. This introduces a dependency on a specific LLM’s behavior. This filter can be readily replaced with more model-agnostic gates. For example, one could use a relevance score threshold from a re-ranker or a lightweight, specialized relevance classifier, similar to the retrieval evaluator in CRAG [65].
@@
D.3 Broader Societal Impact
@@ E.1.3 Attack Generality
As the implemented attacks are generic, their effectiveness may vary across different LLMs. Newer models, such as GPT-4o-mini used in our experiments, incorporate built-in defense mechanisms specifically designed to improve resistance against attacks such as prompt injections [49]. While our attacks proved effective to varying degrees in the evaluation, the presence of such defenses could explain differences in vulnerability observed across models.

### v2

@@ D.1.2 Generality of Weighted Approaches
  • In keyword aggregation in Xiang et al. (2024a), instead of simple counts, one could accumulate the sum of weights of documents supporting each keyword. The filtering threshold $\mu$ could then be applied to these weighted sums.
  • In decoding aggregation in Xiang et al. (2024a), the averaging of next-token probability vectors could become a weighted average, using weights derived from the documents supporting each prediction $v_{j}$.
@@
D.2 Running Time Analysis
While any added latency requires justification, it is crucial to consider the context of modern, potentially complex RAG workflows. Simple RAG involves retrieval and a single generation step, but achieving high quality often necessitates more elaborate strategies. Users interacting with sophisticated RAG systems might experience multi-second latencies, which can stem not only from retrieval and basic generation Milvus (2024); Nguyen et al. (2025) but also from extensive downstream processing, such as reasoning or other test-time scaling techniques applied for enhanced analysis and answer quality.
Our MIS-based approach functions primarily as a document filtering and selection mechanism upstream of this final, potentially costly, answer generation or analysis stage. This contrasts fundamentally with methods such as Keyword Aggregation or Decoding Aggregation Xiang et al. (2024a), which act as alternative inference procedures themselves. The key advantage of our filtering approach is its modularity; it can be seamlessly integrated upstream of any subsequent inference strategy, even though Section 3 presents a specific way that Vanilla RAG is invoked after document selection.
D.3 Limitations and Future Work
@@
Dependency on NLI Model Performance. The efficacy of our MIS-based approach is intrinsically linked to the NLI model’s ability to accurately detect contradiction. Although Theorem 1 accounts for imperfect NLI, the practical impact of more severe NLI inaccuracies, or NLI models that are themselves targeted by sophisticated adversarial examples, deserves more study.
@@
Computational Cost. Although exact MIS is practical for the typical number of retrieved documents (e.g. $k\leq 20$), and our weighted sample and aggregate framework extends scalability, the “isolated answering” step for contradiction graph construction (Section 3.1) does add nontrivial computational latency. While we have demonstrated that this overhead is manageable and have also provided practical speed-up tips, it is a factor to consider. Exploring more efficient methods for contradiction detection can be an interesting future direction.
@@
Heuristic Choices of Parameters and Algorithmic Designs. Our proposed framework incorporates several design choices and parameter settings. For example, Algorithm 1 selects the MIS with the smallest lexicographic order. In our evaluations, we focused on specific configurations such as $m=2$, $T=20$, and using exponentially decaying weights with $\gamma=0.9$. While these configurations have demonstrated strong performance, and Appendix C.6 provides some analysis of how certain parameter choices affect performance, many other reasonable design choices remain interesting to explore. For instance, exploring the use of a maximum weighted independent set could offer a more direct integration of cardinal reliability scores into the MIS selection process itself. Another promising heuristic worth investigating involves applying Algorithm 1 recursively to filter each sampled document set $\mathcal{S}_{t}$ prior to generating intermediate answers in Algorithm 2 (Line 4), which might further bolster the reliability of the final aggregated response.
@@
Reliance on LLM-as-a-Judge. Our empirical evaluations rely on GPT-4o as an LLM-judge for answer correctness and quality. While a common practice, LLM-based evaluation may have inherent biases and may not fully capture all nuances of human assessment.
@@
Exploration of Diverse Adaptive Attack Strategies. Our current work evaluates robustness against several attack types, including corpus poisoning attack, prompt injection attack, and a specific adaptive attack scenario (as detailed in Appendix C.8). However, the landscape of adversarial tactics is continually evolving. To more comprehensively ascertain the resilience of ReliabilityRAG, future work should explore a wider array of sophisticated adaptive attacks. Adversaries with deeper knowledge of the defense mechanism might devise strategies not covered in our present evaluations. A thorough investigation of such advanced adaptive threats would further solidify the understanding of our method’s robustness boundaries and is a valuable direction for continued research.
@@
Scope of Evaluation Benchmarks While our empirical evaluations utilize established datasets such as RQA, NQ, TQA, and the Biography generation dataset, which are common benchmarks in RAG research, it is important to acknowledge a potential limitation shared across much of the current literature. The characteristics and complexities of queries and documents encountered in these datasets may not fully encapsulate the diverse and dynamic nature of real-world web search combined with RAG systems. Consequently, while our results demonstrate significant robustness and utility, performance in live, large-scale commercial search + RAG environments might present additional, unforeseen challenges. This gap between academic benchmarks and real-world deployment scenarios is a broader issue faced by the research community.
@@
Our implementation uses an “I don’t know” response from an LLM to filter irrelevant documents. This introduces a dependency on a specific LLM’s behavior. This filter can be readily replaced with more model-agnostic gates. For example, one could use a relevance score threshold from a re-ranker or a lightweight, specialized relevance classifier, similar to the retrieval evaluator in CRAG Yan et al. (2024).
@@
D.4 Broader Societal Impact
@@ E.1.3 Attack Generality
As the implemented attacks are generic, their effectiveness may vary across different LLMs. Newer models, such as GPT-4o-mini used in our experiments, incorporate built-in defense mechanisms specifically designed to improve resistance against attacks such as prompt injections OpenAI (2024a). While our attacks proved effective to varying degrees in the evaluation, the presence of such defenses could explain differences in vulnerability observed across models.